Design a single equal-angle tension member. It is 4m long and subjected to a tensile
load of 25T. It is connected to a gusseted plate through one leg only. Assume bolt of
16mm. σt = 1500kg/cm2. Check for slenderness ratio. Given rmin = 1.94cm.
Answer #1

solution Given data hood -255 Bole diamerer = 16mm T = 1500 kg lam? Smin - liqucm. equal angle Tension Member Length = um T=Tension applied Permissible Effective area - 25000 kg stress - 1500 kg linh – 25000 con? isom 16.66mm? 1666 mm² = Aett - area1) Take section Another equal angle 150X150X10 A = (150 - 1915–s)x10 - 1275 min A = (150-5)x10 - 1450 min K = 0.725 Aeft = 23